OpenStudy (anonymous):

CAN SOMEONE PLZ HELP?!?!?!?(MEADL+FAN) Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach at high tide measured where 1995 is represented by year 0. Year number Western Beach width (in feet) Dunes Beach width (in feet) 0 100 20 5 90 45 10 80 70 11 78 75 12 76 80 15 70 95 Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table

OpenStudy (anonymous):

@Nasa7 CAN U HELP

OpenStudy (anonymous):

So, The western Beach is losing sand, dunes beach is gaining sand. In a span of about 11 to 12 years

OpenStudy (anonymous):

oh i was over thinking it thanks a ton ur a life saver ive been on this question 4 hrs

OpenStudy (anonymous):

is that all i have to put

OpenStudy (anonymous):

No I am still typing :P

OpenStudy (anonymous):

I am beginning to explain it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, the western beach looses 2 ft while dunes beach get 5ft more per year.

OpenStudy (anonymous):

So, we would get a constant slope when you plot the graphs so 2 less per year for western and 5 more per year for dunes so both would be linear

OpenStudy (anonymous):

ok i get it thanks for explaining ,i dont think i have to graph it just write it out

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

can u check my work for Between which years will the beaches have approximately the same width? 11 and 12/2006-2007 Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width -2x + 100 = 5x + 20 100 - 20 = 7x 80 = 7x 80/7 = x x ~11.43 years

OpenStudy (anonymous):

So what I am going to have you do is there are many posts just like this read through this post then tell me if you have any other questions :)

OpenStudy (anonymous):

ok thanks for helping ; )

OpenStudy (anonymous):

No problem thanks for the medal :)

OpenStudy (triciaal):

I answered this question before too. including a sketch of the functions.

OpenStudy (anonymous):

Ah yes also take a look at triciaal's let me get it :)

OpenStudy (anonymous):

ok thanks again

OpenStudy (anonymous):

Thank you @triciaal

OpenStudy (triciaal):

welcome Is there a way to quickly retrieve the original ( this was phi's) or in general to search closed questions?